Let matrix M =

-8 -24 12
0 4 0
6 12 10

(a) Find the eigenvalues of M

(b) For each eigenvalue λ of M, find a basis for the eigenspace of λ.

(c) Is the matrix M diagonalizable? If so, find matrices D and P such that D is a diagonal matrix and M=PDP−1. If not, explain carefully why not.
Answer #1

Ansi M - - 24 4. - 12 O 4 % 12 10 M-d=0 -24 -12 -8-d O 6 0 4-d 12 Ji lord -12 14-21 (4-211-8-2 6 lord 2-6 (4-2) [(x-10) (8+d)Ex=4= {[:(312 5 . Corresponding = -2 (M+21)x=0 .6 - 24 o O 6 X2 23 6 12. X2=0 x + 4x + 3 = 0 = X, 2x3 = 0 Ed=-2- 1 -1/2 therethumbs up - 12 13. N P = o O O - 9+3JTI 9 +355 (C) M is diagonalizable because there is matrix inventible mat aux p such that M= PDPT D thumbs up 12 6 6 12 10 O 6 10-d 12 Now we will find eigen vectors. Ans) M = -8 - 2 4 O 4. а) IM-XI = 0 1-8-d -24 12. 4-d 6 (4-0) [ 68-1 Ex,= 1:35 is eigen vector 2 for 12 = 1-3557. - 24 12 M [-g +3577 3+37 22 6 12 3 9+3 T X₂=0 1-9 + 3877) X, + 12 X₂ = 0 Xi -)2.